OpenStudy (anonymous):

I really don't understand this question. Use complete sentences to describe a real-world scenario that could be represented by the inequality 4x + 2y 40. Part 2: Choose one ordered pair that is a solution to the given inequality and explain what that ordered pair means in the context of your real-world scenario.

OpenStudy (anonymous):

you sell beer for $4 and soda for $2. if x is number of beers and y is number of sodas, how many do you have to sell to make at least $40

OpenStudy (anonymous):

Part 2: With above soda/beer scenario: 4x+2y=40 We can just choose a value for x or y Let's say we sell 2 beers, then we can substitute x with 2 4(2)+2y=40...and solve 8+2y=40 2y=32 y=16, therefore selling 2 beers and 16 sodas brings in $40